How much would you loan your sister if he said he could repay you $100 in six months, $200 in a year, and $500 in two years if you can get 2 percent interest from the bank on a six month CD? Show that you are in fact indifferent between the loan and putting your money in the bank for the next two years.
Answer #1
1) Amount that can be loaned = 100/1.02+200/1.02^2+500/1.02^4 = $             752.20
2) FV if invested in CD would be 752.20*1.02^4 = $             814.21
FV if the repayments are invested = 100*1.02^3+200*1.02^2+500 = $             814.20
AS the FVs of both the options are equal one should be indifferent
as to the two alternatives of giving loan to sister or investing in a CD.